# Qimax: Efficient quantum simulation via GPU-accelerated extended stabilizer formalism

### v1

@@ 1 Introduction
Simulating a quantum system classically demands significant resources, scaling exponentially with the number of particles, posing a challenging computational problem. For instance, Google’s quantum computer, Sycamore, demonstrated quantum supremacy by producing results in days that would take classical computers thousands of years to simulate [1]. Various simulation methods have been proposed, including state-vector [2], tensor-network [3], decision diagram [4], and stabilizer formalism [5]. Each method offers different levels of fidelity and resource requirements depending on the quantum circuit’s size and type.
@@
Stabilizer formalism based on Heisenberg’s picture provides a compact representation for quantum systems and gate applications where the quantum system can be described as stabilizers. Quantum gates act on the stabilizers by mapping between $n$-qubit Pauli string $\mathcal{P}_{n}$ rather than matrix-vector multiplication. Existing stabilizer formalism packages have been developed, although they can run a hundred-qubit Clifford circuit with a million gates, the stabilizer formalism limits the application of the quantum simulator, especially for algorithms that require non-Clifford gates. On the other hand, using non-Clifford gates in (extended) stabilizer formalism may lead to a fast increase in complexity. For instant, [6]shows the exponential rise of stabilizer rank in cases of QNN, which require non-Clifford gates. Other studies address this by employing approximation methods, balancing complexity and accuracy [7].
There are two use cases of stabilizer formalism: (1) large-qubit Clifford circuits, and (2) low-rank Clifford and non-Clifford. Because Stim [8] effectively handles the first, our research focuses on the second. Clifford circuits are only used for quantum error correction and benchmarking the performance of quantum hardware. Near-Clifford circuits have wider applications. Table 1 summarizes recent stabilizer and extended stabilizer formalism packages, revealing that all prior extended stabilizer packages rely on sequential algorithms, limiting their performance on GPUs. Our proposed package, Qimax, addresses this limitation by offering three modes: v1 (standard extended stabilizer formalism, CPU-based), v2 (parallel stabilizer formalism, GPU-based, using one-hot vector Pauli encoding), and v3 (parallel stabilizer formalism, GPU-based, using tensor index encoding).
@@
stim [8] Python C++ Sequential Clifford 2021
PyClifford Python Python Sequential Clifford Qiskit [2], Pyscf 2022
QuiZX [9] Python Rust Sequential Clifford + non Clifford ZX Calculus [10] 2022
WMC [6] Python Python Sequential Clifford + non Clifford GPMC solver, d4max solver 2024
QuaSARQ [11] C++ C++ (CUDA) Parallel Clifford 2025
Qimax Python Python Sequential or Parallel Clifford + non Clifford CuPy [12] 2025
@@ 2.1 Extended stabilizer formalism
An $n$-qubit quantum state $|\psi\rangle$ is defined as a stabilizer state by an operator $U$ if and only if it is a $+1$ eigenvector of $U$, meaning $U|\psi\rangle=|\psi\rangle$. The stabilizer states form a strict subset of all quantum states that can be uniquely described by the stabilizer group $\{\mathcal{P}_{n}\}$. The elements of the stabilizer group are called stabilizers. Any stabilizer group can be specified by a set of generators $G_{|\psi\rangle}\equiv G=\{\mathbb{P}_{n}\}$ so that every element in the stabilizer group can be obtained through matrix multiplication between any pair $\{\mathbb{P}_{n,i},\mathbb{P}_{n,j}\}$. The number of generators (order) of $G$ is always $n$. These generators evolve under the application of $m$ quantum gates, as shown in Equation (7):
@@
Example 1
Consider a 3-qubit circuit $CX_{0,1}R_{z}(\pi/3)_{0}SX_{0}|000\rangle$. he evolution of the stabilizer generators $G$ is tracked gate by gate, as specified in Table 2. For instance, the Pauli string $-Y$ is converted to $(\frac{1}{2}X-\frac{\sqrt{3}}{2}Y)$:
$\left\langle\begin{array}[]{c}Z_{3,0}\\ Z_{3,1}\\ Z_{3,2}\\ \end{array}\right\rangle\xrightarrow{SX_{0}}\left\langle\begin{array}[]{c}-YII\\ IZI\\ IIZ\\ \end{array}\right\rangle\xrightarrow{R_{z}(\frac{\pi}{3})_{0}}\left\langle\begin{array}[]{c}\frac{1}{2}XII-\frac{\sqrt{3}}{2}YII\\ IZI\\ IIZ\\ \end{array}\right\rangle\xrightarrow{CX_{0,1}}\left\langle\begin{array}[]{c}\frac{1}{2}XXI-\frac{\sqrt{3}}{2}YXI\\ ZZI\\ IIZ\\ \end{array}\right\rangle$ (9)
We can relate $G^{(m)}$ directly to the target density operator either as a product of stabilizers or as a sum of Pauli operators, as shown in (10) and (11), respectively:
$\rho^{(m)}\equiv|\psi^{(m)}\rangle\langle\psi^{(m)}|$ $=\frac{1}{2^{n}}\prod_{j=1}^{n}\left(I^{\otimes n}+\mathbb{P}^{(m)}_{n,j}\right)$ (10)
$=\frac{1}{2^{n}}\sum_{\mathcal{P}_{n}\in\mathcal{P}_{|\psi\rangle}}\lambda_{\mathcal{P}_{n}}\mathcal{P}_{n}.$ (11)
@@ 3.1 Instruction
We present gates as instructions similar to Qiskit [2]. Each instruction is a tuple $\{g,w,\theta\}$, representing the gate name, qubit (wire), and parameter value (if applicable), respectively. For gate without parameters, $\theta$ is set as 00. Performing gate-by-gate on stabilizers leads to low performance; hence, we consider performing operators instead, where each operator contains gates of the same type. A quantum circuit is partitioned into $K$ $U_{k}$ (only 1-qubit gates) and $K^{\prime}$ $V_{k}$ (only 2-qubit gates) because the 2-qubit gate has different behavior from other gates. These operators interleaved each other, as presented in Fig. 2, obey $|K-K^{\prime}|\leq 1$. The operator sequence is determined by the first instruction: if it is a single-qubit gate, the order will be $U_{0},V_{1},U_{1},V_{1},\ldots$; otherwise, it starts with $V_{0}$. Within each $U_{k}$, gates are further grouped by qubit, with $U_{k,j}$ denoting instructions in the $k^{\text{th}}$ single-qubit operator acting on the $j^{\text{th}}$ qubit. By grouping gates, there is no difference between Clifford and non-Clifford gates.
@@ 3.2 Encoding
We encode any Pauli string $P_{n}$ as an integer index using base-4 with the mapping $I=0,X=1,Y=2,Z=3$. The index ranges from 00 (representing $I^{\otimes n}$) to $14^{n}-$ 1 ($Z^{\otimes n}$). Conversely, decoding an index back to a Pauli string requires #Qubits. When a non-Clifford gate acts on a Pauli matrix $p\neq I$ in $P_{n}$, it may transform into a linear combination $p=w_{0}X+w_{1}Y+w_{2}Z\equiv[w_{0},w_{1},w_{2}]$ as illustrated in Example. 9. We call $w_{j}$ the weight of transformed generators. Since every Pauli matrix $p_{j}$ in a generator may take this form, applying non-Clifford gates can transform a generator from a simple form $\mathbb{P}_{n}=\sum_{j}\lambda_{j}P_{n,j}$ (a sum of Pauli strings) to complex form ^$\hat{\mathbb{P}}_{n}=\sum_{j}\lambda_{j}\prod_{k=1}^{n}\left(\sum_{l}w_{k,l}p_{k.l}\right)$, where the product is over qubits and the sum is over Pauli matrices, as shown in Example. 3.
@@ 3.3 Gate
for $k$ in $[0\ldots K-1]$ in [parallel] do
     for $j$ in $[0\ldots n-1]$ in [parallel] do
@@
For simply, we consider the CX gate, which acts on a control qubit (indexed by $c$) and a target qubit (indexed by $t$), modifying the $c^{\text{th}}$ and $t^{\text{th}}$ Pauli matrices in each Pauli string $P_{n,j}$. The CX gate transforms these Pauli matrices according to specific rules, such as: $XI\leftrightarrow XX,IY\leftrightarrow ZY,YI\leftrightarrow YX,IZ\leftrightarrow ZZ,XY\leftrightarrow YZ$ while leaving $II,IX,ZI$ and $ZX$ unchanged, or change both Pauli and sign $(XZ\leftrightarrow-YY)$. For $m^{\prime}$ CX gates, each Pauli string is updated according to these rules $m^{\prime}$ times. These transformations are efficiently implemented using a lookup table, as shown below:
$\text{LUT}_{c}=\begin{bmatrix}0&0&3&3\\ 1&1&2&2\\ 2&2&1&1\\ 3&3&0&0\end{bmatrix},\text{LUT}_{t}=\begin{bmatrix}0&1&2&3\\ 1&0&3&2\\ 1&0&3&2\\ 0&1&2&3\end{bmatrix},$
$\text{LUT}_{\text{sign}}=\begin{bmatrix}1&1&1&1\\ 1&1&1&-1\\ 1&1&-1&1\\ 1&1&1&1\end{bmatrix}.$
@@ 5 Implementation
[…]
@@ 6 Experiments
[…]
@@
Qimax v1 shares properties with other stabilizer formalism packages, such as Stim [8], enabling the simulation of many-qubit Clifford circuits in seconds, as shown in Figure4, in this cases, GHZ and Graph circuit has the average stabilizer rank as 1. However, Qimax v1 is slower than Stim due to its support for non-Clifford gates, prioritizing functionality over performance. Qimax v3 demonstrates superior performance in experiments with large #Repeats, as shown in Figure5, outperforming GPU-accelerated Qiskit and Pennylane. At higher #Repeats, Qimax v3’s performance advantage over these packages increases significantly. CPU-based versions of all packages exhibit the longest execution times, with Qiskit generally outperforming Pennylane. All packages exhibit exponential scaling with #Qubits due to the state-vector size (for Qiskit and Pennylane) or the increase in gate count and average stabilizer rank (for Qimax).
Figure 4: The execution time for GHZ circuit and graph circuit on Qimax, these circuits are generated from MQT Bench [14].
@@ 7 Conclusion
Acknowledgment
This work was supported by JST-ALCA-Next Program Grant Number JPMJAN23F4, Japan, and the Next Generation Researchers Challenging Research Program under number zk24010019. The research has been partly executed in response to the support of JSPS, KAKENHI Grant No. 22H00515, Japan.
@@ Appendix A Utilities function used in three mode
  • $\mathbb{P}$) and index_to_word($\mathbb{P},n$): Encode a Pauli word as integer using base-4. On the other hand, we need an additional parameter #Qubits.
  • index_to_indices($i$) and indices_to_index($I$): Convert an encoded Pauli word (integer form) to array form and vice versa. For example: $27(XYZ)\rightarrow[1,2,3]$.
[…]
  • create_chain ($K,K^{\prime}$): Create the chain $[1,0,1,\ldots]$ if the instructions begin with a 1-qubit gate and vice versa.
@@ Appendix B Data for constructing 1-qubit lookup table
$X$ $Z$ $-Z$ $[w_{0},w_{1},-w_{2}]$
$SX$ $Y$ $Z$ $[w_{0},-w_{2},w_{1}]$
$Z$ $-Y$
$R_{x}(\theta)$ $Y$ $\cos(\frac{\theta}{2})Y+\sin(\frac{\theta}{2})Z$ $[w_{0},w_{1}\cos(\theta)-w_{2}\sin(\theta),w_{1}\sin(\theta)+w_{2}\cos(\theta)]$
@@
$R_{y}(\theta)$ $X$ $\cos(\frac{\theta}{2})X-\sin(\frac{\theta}{2})Z$ $[w_{0}\cos(\theta)+w_{2}\sin(\theta),w_{1},w_{2}\cos(\theta)-w_{0}\sin(\theta)]$
@@
$R_{z}(\theta)$ $X$ $\cos(\frac{\theta}{2})X+\sin(\frac{\theta}{2})Y$ $[w_{0}\cos(\theta)-w_{1}\sin(\theta),w_{1}\cos(\theta)+w_{0}\sin(\theta),w_{2}]$
@@
Table 2: Quantum gate operations and results with $\mathbf{w}^{(t)}=[w_{0},w_{1},w_{2}]$ on $\{H,S,X,SX,R_{x}(\theta),R_{y}(\theta),R_{z}(\theta)\}$. The unchanged cases are ignored.

### v2

@@ Abstract
keywords:
quantum simulation, stabilizer, CUDA, Cupy
@@ 1 Introduction
Simulating a quantum system classically demands significant resources, scaling exponentially with the number of particles, posing a challenging computational problem. For instance, Google’s quantum computer, Sycamore, demonstrated quantum supremacy by producing results in days that would take classical computers thousands of years to simulate [sycamore]. Various simulation methods have been proposed, including state-vector [javadiabhari2024quantumcomputingqiskit], tensor-network [10313722], decision diagram [Vinkhuijzen2023limdddecision], and stabilizer formalism [Gidney2021stimfaststabilizer]. Each method offers different levels of fidelity and resource requirements depending on the quantum circuit’s size and type.
@@
Stabilizer formalism based on Heisenberg’s picture provides a compact representation for quantum systems and gate applications where the quantum system can be described as stabilizers. Quantum gates act on the stabilizers by mapping between $n$-qubit Pauli string $\mathcal{P}_{n}$ rather than matrix-vector multiplication. Existing stabilizer formalism packages have been developed, although they can run a hundred-qubit Clifford circuit with a million gates, the stabilizer formalism limits the application of the quantum simulator, especially for algorithms that require non-Clifford gates. On the other hand, using non-Clifford gates in (extended) stabilizer formalism may lead to a fast increase in complexity. For instant, [modelcounting]shows the exponential rise of stabilizer rank in cases of QNN, which require non-Clifford gates. Other studies address this by employing approximation methods, balancing complexity and accuracy [Bravyi2019simulationofquantum].
There are two use cases of stabilizer formalism: (1) large-qubit Clifford circuits, and (2) low-rank Clifford and non-Clifford. Because Stim [stim] effectively handles the first, our research focuses on the second. Clifford circuits are only used for quantum error correction and benchmarking the performance of quantum hardware. Near-Clifford circuits have wider applications. Table 1 summarizes recent stabilizer and extended stabilizer formalism packages, revealing that all prior extended stabilizer packages rely on sequential algorithms, limiting their performance on GPUs. Our proposed package, Qimax, addresses this limitation by offering three modes: v1 (standard extended stabilizer formalism, CPU-based), v2 (parallel stabilizer formalism, GPU-based, using one-hot vector Pauli encoding), and v3 (parallel stabilizer formalism, GPU-based, using tensor index encoding).
@@
stim [stim] Python C++ Sequential Clifford 2021
PyClifford Python Python Sequential Clifford Qiskit [javadiabhari2024quantumcomputingqiskit], Pyscf 2022
QuiZX [zxpackage] Python Rust Sequential Clifford + non Clifford ZX Calculus [kissinger2020Pyzx] 2022
WMC [modelcounting] Python Python Sequential Clifford + non Clifford GPMC solver, d4max solver 2024
QuaSARQ [gpu_stabilizer] C++ C++ (CUDA) Parallel Clifford 2025
Qimax Python Python Sequential or Parallel Clifford + non Clifford CuPy [cupy_learningsys2017] 2025
@@ 2.1 Extended stabilizer formalism
An $n$-qubit quantum state $|\psi\rangle$ is defined as a stabilizer state by an operator $U$ if and only if it is a +1+1 eigenvector of $U$, meaning $U|\psi\rangle=|\psi\rangle$. The stabilizer states form a strict subset of all quantum states that can be uniquely described by the stabilizer group $\{\mathcal{P}_{n}\}$. The elements of the stabilizer group are called stabilizers. Any stabilizer group can be specified by a set of generators $G_{|\psi\rangle}\equiv G=\{\mathbb{P}_{n}\}$ so that every element in the stabilizer group can be obtained through matrix multiplication between any pair $\{\mathbb{P}_{n,i},\mathbb{P}_{n,j}\}$. The number of generators (order) of $G$ is always $n$. These generators evolve under the application of $m$ quantum gates, as shown in Equation (7):
@@
Example 1.
Consider a 3-qubit circuit $CX_{0,1}R_{z}(\pi/3)_{0}SX_{0}|000\rangle$. he evolution of the stabilizer generators $G$ is tracked gate by gate, as specified in Table 2. For instance, the Pauli string Y-Y is converted to $(\frac{1}{2}X-\frac{\sqrt{3}}{2}Y)$:
[…]
We can relate $G^{(m)}$ directly to the target density operator either as a product of stabilizers or as a sum of Pauli operators, as shown in (10), respectively:
$\rho^{(m)}\equiv|\psi^{(m)}\rangle\langle\psi^{(m)}|=\frac{1}{2^{n}}\prod_{j=1}^{n}\left(I^{\otimes n}+\mathbb{P}^{(m)}_{n,j}\right)=\frac{1}{2^{n}}\sum_{\mathcal{P}_{n}\in\mathcal{P}_{|\psi\rangle}}\lambda_{\mathcal{P}_{n}}\mathcal{P}_{n}.$ (10)
@@ 3.1 Instruction
We present gates as instructions similar to Qiskit [javadiabhari2024quantumcomputingqiskit]. Each instruction is a tuple $\{g,w,\theta\}$, representing the gate name, qubit (wire), and parameter value (if applicable), respectively. For gate without parameters, $\theta$ is set as 0. Performing gate-by-gate on stabilizers leads to low performance; hence, we consider performing operators instead, where each operator contains gates of the same type. A quantum circuit is partitioned into $K$ $U_{k}$ (only 1-qubit gates) and $K^{\prime}$ $V_{k}$ (only 2-qubit gates) because the 2-qubit gate has different behavior from other gates. These operators interleaved each other, as presented in Fig. 2, obey $|K-K^{\prime}|\leq 1$. The operator sequence is determined by the first instruction: if it is a single-qubit gate, the order will be $U_{0},V_{1},U_{1},V_{1},\ldots$; otherwise, it starts with $V_{0}$. Within each $U_{k}$, gates are further grouped by qubit, with $U_{k,j}$ denoting instructions in the $k^{\text{th}}$ single-qubit operator acting on the $j^{\text{th}}$ qubit. By grouping gates, there is no difference between Clifford and non-Clifford gates.
@@ 3.2 Encoding
We encode any Pauli string $P_{n}$ as an integer index using base-4 with the mapping $I=0,X=1,Y=2,Z=3$. The index ranges from 0 (representing $I^{\otimes n}$) to $4^{n}-1$ ($Z^{\otimes n}$). Conversely, decoding an index back to a Pauli string requires #Qubits. When a non-Clifford gate acts on a Pauli matrix $p\neq I$ in $P_{n}$, it may transform into a linear combination $p=w_{0}X+w_{1}Y+w_{2}Z\equiv[w_{0},w_{1},w_{2}]$ as illustrated in Example. 9. We call $w_{j}$ the weight of transformed generators. Since every Pauli matrix $p_{j}$ in a generator may take this form, applying non-Clifford gates can transform a generator from a simple form $\mathbb{P}_{n}=\sum_{j}\lambda_{j}P_{n,j}$ (a sum of Pauli strings) to complex form $\hat{\mathbb{P}}_{n}=\sum_{j}\lambda_{j}\prod_{k=1}^{n}\left(\sum_{l}w_{k,l}p_{k.l}\right)$, where the product is over qubits and the sum is over Pauli matrices, as shown in Example. 3.
@@ 3.3 Gate
for $k$ in $0\ldots K-1]$ in [parallel] do
  for $j$ in $0\ldots n-1]$ in [parallel] do
@@
For simply, we consider the CX gate, which acts on a control qubit (indexed by $c$) and a target qubit (indexed by $t$), modifying the $c^{\text{th}}$ and $t^{\text{th}}$ Pauli matrices in each Pauli string $P_{n,j}$. The CX gate transforms these Pauli matrices according to specific rules, such as: $XI\leftrightarrow XX,IY\leftrightarrow ZY,YI\leftrightarrow YX,IZ\leftrightarrow ZZ,XY\leftrightarrow YZ$ while leaving II,IX,ZIII,IX,ZI and $ZX$ unchanged, or change both Pauli and sign $(XZ\leftrightarrow-YY)$. For $m^{\prime}$ CX gates, each Pauli string is updated according to these rules $m^{\prime}$ times. These transformations are efficiently implemented using a lookup table, as shown below:
$\text{LUT}_{c}=\begin{bmatrix}0&0&3&3\\ 1&1&2&2\\ 2&2&1&1\\ 3&3&0&0\end{bmatrix},\text{LUT}_{t}=\begin{bmatrix}0&1&2&3\\ 1&0&3&2\\ 1&0&3&2\\ 0&1&2&3\end{bmatrix},\text{LUT}_{\text{sign}}=\begin{bmatrix}1&1&1&1\\ 1&1&1&-1\\ 1&1&-1&1\\ 1&1&1&1\end{bmatrix}$. (11)
@@ 5 Implementation
The three modes of Qimax are implemented in Python 3.10.11, using CuPy-cuda12x 13.4.0 [cupy_learningsys2017] as the backend for GPU acceleration. Computationally intensive functions are executed in CUDA kernels with a block size of 256 and a grid size of $\lceil(\text{total elements}+255)/256\rceil$, where total elements represent the number of tensor entries or tasks. The code and data for this study are accessible at https://github.com/NAIST-Archlab/qimax-soft.
@@ 6 Experiments
Comparable software includes Pennylane 0.41.0 [bergholm2022pennylaneautomaticdifferentiationhybrid] and Qiskit 1.1.1 [javadiabhari2024quantumcomputingqiskit], both accelerated with NVIDIA’s cuQuantum SDK (lightning.gpu 0.41.0 and cusvaer_simulator_statevector 0.13.3, respectively). Experiments were conducted on an Intel i9-10940X CPU (3.30 GHz) and an NVIDIA RTX 4090 GPU. Execution time was measured from the initialization of stabilizer generators to the computation of final generators. Each experiment was run at least 10 times, and the average execution time was calculated. The benchmarked circuits include the GHZ circuit (multi-qubit entangled state), Graph circuit (graph-based quantum state), and $|XYZ+chain\rangle$, as shown in Figure2 (b). The circuit structure consists of $L$ layers, with the XYZ component repeated $\text{Repeats}$ times within each layer. Experiments varied $\text{Repeats}$ from 100 to 100,000 to evaluate Qimax’s scalability. The #Gates increases linearly with $\text{Repeats}$ and #Qubits. The average stabilizer rank of generators, which impacts performance, is plotted in Figure3; higher ranks correspond to lower performance in Qimax due to increased computational complexity.
@@
Qimax v1 shares properties with other stabilizer formalism packages, such as Stim [stim], enabling the simulation of many-qubit Clifford circuits in seconds, as shown in Figure4, in this cases, GHZ and Graph circuit has the average stabilizer rank as 1. However, Qimax v1 is slower than Stim due to its support for non-Clifford gates, prioritizing functionality over performance. Qimax v3 demonstrates superior performance in experiments with large #Repeats, as shown in Figure5, outperforming GPU-accelerated Qiskit and Pennylane. At higher #Repeats, Qimax v3’s performance advantage over these packages increases significantly. CPU-based versions of all packages exhibit the longest execution times, with Qiskit generally outperforming Pennylane. All packages exhibit exponential scaling with #Qubits due to the state-vector size (for Qiskit and Pennylane) or the increase in gate count and average stabilizer rank (for Qimax).
Figure 4: The execution time for GHZ circuit and graph circuit on Qimax, these circuits are generated from MQT Bench [quetschlich2023mqtbench].
@@ 7 Conclusion
[…]
Acknowledgements
This research was supported by the VNUHCM - University of Information Technology’s Scientific Research Support Fund.
@@ Appendix A Utilities function used in three mode
  • word_to_index($\mathbb{P}$) and index_to_word($\mathbb{P},n$): Encode a Pauli word as integer using base-4. On the other hand, we need an additional parameter #Qubits.
  • index_to_indices(ii) and indices_to_index($I$): Convert an encoded Pauli word (integer form) to array form and vice versa. For example: $27(XYZ)\rightarrow[1,2,3]$.
  • $\text{create\_Z}_{j}(j,n)$: Return $n$-dim one-hot vector $\bm{e}_{j}$.
  • create_chain ($K,K^{\prime}$): Create the chain $1,0,1,\ldots]$ if the instructions begin with a 1-qubit gate and vice versa.
@@ Appendix B Data for constructing 1-qubit lookup table
$X$ $Z$ Z-Z $w_{0},w_{1},-w_{2}]$
SXSX $Y$ $Z$ $w_{0},-w_{2},w_{1}]$
$Z$ Y-Y
$R_{x}(\theta)$ $Y$ $\cos(\frac{\theta}{2})Y+\sin(\frac{\theta}{2})Z$ $w_{0},w_{1}\cos(\theta)-w_{2}\sin(\theta),w_{1}\sin(\theta)+w_{2}\cos(\theta)]$
@@
$R_{y}(\theta)$ $X$ $\cos(\frac{\theta}{2})X-\sin(\frac{\theta}{2})Z$ $w_{0}\cos(\theta)+w_{2}\sin(\theta),w_{1},w_{2}\cos(\theta)-w_{0}\sin(\theta)]$
@@
$R_{z}(\theta)$ $X$ $\cos(\frac{\theta}{2})X+\sin(\frac{\theta}{2})Y$ $w_{0}\cos(\theta)-w_{1}\sin(\theta),w_{1}\cos(\theta)+w_{0}\sin(\theta),w_{2}]$
@@
Table 2: Quantum gate operations and results with $w^{(t)}=[w_{0},w_{1},w_{2}]$ on $\{H,S,X,SX,R_{x}(\theta),R_{y}(\theta),R_{z}(\theta)\}$. The unchanged cases are ignored.